Given the following code…

var num1 = 31;

var num2 = 96;

var num3 = 22;

var temp = num2;

num3 = temp + num1;

num1 = num1 + num3

var num4 = num3 + 11;

num2 = num2 + num4;

var ans = num1 + num2 + num3;

What is the final value that is assigned to ‘ans’?
Answer #1

Answer: Hey dear student finds the solution of your query if you have any doubt feel free to ask. Thanks!!

final value ans = 519

let's see: put value in place of variables

num1 = 31

num2 = 96

num3 = 22

temp = 96 (num2)

num3 = 96 + 31 = 127 (temp + num1)

num1 = 31 + 127 = 158 (num1 + num3)

num4 = 127 + 11 = 138 (num3 + 11)

num2 = 96 + 138 = 234 (num2 + num4)

ans = 158 + 234 +127 = 519 (num1 + num2 + num3)

Hence we got 519 in ans variable. If any confusion comment me in comment box.